A Complete Bibliography of ACM Transactions on Economics and Computation (TEAC)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

14 June 2024 Version 1.33

Title word cross-reference

α [CR17]. ϵ [Elb21]. ∃ [GMVY18]. ∃R [BM21]. λ [GPZ21].

-complete [BM21]. -Envy-Free [Elb21]. -regular [GPZ21]. -Strongly [CR17].

1 [CH21a, HCGM22]. 1/e [ST20].

2 [CH21b].

Acceptance [Mau21]. ACM [CM13, CM17, GÜ23]. Ad [CV15, NW18]. Additive [GJLZ20]. Adversarial [EKM18]. Advertising [CMV15, GMMV15, GMS15, NW18]. Affine [NS15]. After [CCS17]. Agenda [ABC^{+}15]. Agents [CCK^{+}16, DSE15, GK16, KR20, MXX22, SY22, YOA^{+}13]. Aggregation [KKL19]. Algorithm [GHV23]. Algorithmic [AER^{+}24, FN13]. Algorithms [BK20, SSCD14]. Allocation [AV20, CISZ21, CEK16, DSSX21, EKM18, GHLTC23, GHBBM22, NW18]. Allocations [BCIZ21, KPS18]. Alternating [CEK16]. Altruism [CDKS14, HS13]. Amazon [BYBYST13]. among [MXX22]. Analysis [KFH15, SSO23, JP13]. Anarchy [AER^{+}24, AV23, BGR14, CDKS14, CDD^{+}14, CKST16, KS23, MBP22, PNS16, Rou15]. Anonymity [RS16]. Answering [SY22]. Ante [Mau21]. Antifolk [PRU17]. Applications [CR17, RW18]. Approach [EG16]. Approval [SZ20]. Approximate [CFGS15, FS16, GR16, PT13]. Approximately [CBGdK^{+}20]. Approximating [AKS17]. Approximation [AZ21, BK20]. Arbitrage [CMV15]. Arrival [DSS^{+}24]. Arrivals [DSS^{+}24]. Arrow [DGV16]. Assignment [BCH^{+}20, DSS^{+}24, FB17, FSZ18]. Assortment [ILM^{+}21]. Atomic [HS13, PCFM21]. Auction [AKLV17, GHV23]. Auctioning [ADH15]. Auctions [AV23, AMS14, Ass20, ADH15, BL16, CDE^{+}14, CMV15, CKST16, CR17, CBLHS15, CBGdK^{+}20, DFI14, DHS15, DHW15, DHS18, HIMM13, HK15, KP13, Ma22, RGKK16]. Augmentations [ISW23]. Automated [OPRS13]. Averse [DSE15]. Axiomatic [FSST19].

Back [CM17]. Bad [EPR19]. Badge [EG16]. Balancing [VBMM22]. Based [BB17, DFJ^{+}15, DHS18, GMS15, JZ23, BCH^{+}20, GHBBM22, GH24]. Bayesian [CR17, GK17]. Be [CKV20, EPR19]. Behavior [FP16]. Behavioral [EPR19]. Best [CDE^{+}14, DS22, GMC21]. Best-Response [CDE^{+}14, GMC21]. Better [HR20]. Beyond [PPS15]. Bid [HN19]. Bid-Limited [HN19]. Bidders [DHK^{+}16, HK15]. Bidding [KMM24, SG20]. Bimatrix [FS16]. Binary [ABC^{+}15, LSA24, YOA^{+}13]. Bitcoin [LSZ22]. Black [GLT20]. Black-Box [GLT20]. Blockchain [GH24]. Blockchain-based [GH24]. Boolean [GNR18]. Border [GNR18]. Borders [GNR18]. Bound [MBP22]. Boundary [JLSW17]. Bounded [BGP15]. Bounded-Distance [BGP15]. Bounds [CKST16, GR16, KS23]. Box [GLT20]. Bribing [KKM20]. Broadcast [CNP^{+}21]. Budget [CGMW21, DDW18, DHS15, GJLZ20, JT21]. Budget-Constrained [CGMW21]. Budget-Feasible [GJLZ20]. Budgeting [GKSA19]. Budgets [CBLHS15]. Buy-in [SSO23]. Buyer [AV23, CGMW21, RW18]. Buyers [Ma22].

Call [WS15]. Can [CKV20, EPR19]. Candidate [AZ21]. Capacity [FG16]. Cardinal [GH18]. Case [ABM20, BGR14, GH24]. Catastrophe [LSCP23]. Causal [IO18]. Characteristic [MSR^{+}14]. Characterization [FSST19]. Characterizing [KLM21]. Checking [HH13]. Chief [GÜ23, PS17]. Choice [AZ21]. Choices [BFK17]. Chores [GHLTC23]. Classifiers [DFJ^{+}15, KR20]. Closing [ST20]. Cloud [KLM21, KKS19, SG20]. Clustering [KS23]. Co [GÜ23, PS17]. Co-Editors-in-Chief [GÜ23, PS17]. Coalition [BBW23]. Coalitions [FG16]. Collective [ABC^{+}15]. Combinatorial [Ass20, BMS22, BL16, CEK16, CBGdK^{+}20, DHS18, ILM^{+}21]. Committee [CJMW20, FSST19]. Common [CCS17]. Communication [GSP21]. Communities [ZPvdS14]. Comparative [HV16]. Competition [PST14]. Competitive [HPS^{+}18, KP13, Sar13]. complete [BM21]. Completeness [GMVY18]. Complexity [AM14, BB15, CR22, GSP21, GLT20, GPS13, GR16, OPR16, SSCD14]. Compressions [DHS18]. Computable [DS22]. Computation [BBH^{+}17, CFGS15, GSP21, HMV16, MSR^{+}14, CM13, GÜ23]. Computational [HP19]. Computing [BB17, GMC21, JLSW17, KT23, SSO23]. Concepts [BB17, JZ23]. Congestion [BGR14, BV19, CFGS15, CG16, GKK20, HS13, KR15, PCFM21]. Connected [AD15]. Connection [ACV13]. Consensus [HH13, MT13]. Constrain [FG16]. Constrained [CGMW21]. Constraints [DDW18, FG16, FSZ18]. Contagious [BEK^{+}13]. Contest [GK16]. Contests [GH18, LSA24]. Continuous [HLPV18].

Convergence [AKLV17, CDE^{+}14]. Convex [ACV13, DGV16, GMC21]. Convexity [KMM24]. Cookie [GMMV15]. Copies [GHLTC23]. Correlated [ABF^{+}17, BB15, MS15]. Correlations [ISW23]. Cost [BMS22, GKK20, RS16, SSO23]. Cost-Sharing [GKK20, RS16]. Costs [KC21]. Covering [BKPS14, MS15]. Creation [AEED^{+}14, BGP15]. Cross [HH13]. Crowdsourcing [SZ20]. cryptographic [GLR13]. Current [CEK16].

Data [GILR20, WYZ18]. Debreu [DGV16]. Decentralized [GH24]. Decision [BM21, CKRS14, FP16, GMVY18]. Decisions [ABC^{+}15]. Deconstructing [BYBYST13]. Deferred [Mau21]. Delay [SG20]. Delivery [CV15]. Demand [LGMC22]. Demands [GHV23]. Democracy [GKMP21]. Design [BFT16, BL16, EG16, FZ15, GLT20, GK16, HH13, HMV16, KNP13, KC21, LSCP23, MS15, PT13, RTC16, XV15]. Designing [KS19]. Destabilize [LSCP23]. Deterministic [FT14, Ma22]. Dichotomous [ABM20]. Differentiated [SG20]. Digital [AMS14]. Discounted [DS22]. Discriminant [DFJ^{+}15]. Discriminant-Based [DFJ^{+}15]. Discrimination [KM24]. Display [CMV15, GMS15]. Distance [BGP15]. Distortion [EKPS24]. Distributed [ADH19, ASA17, FTW20, JLSW17]. Distribution [JZ23]. Distribution-Based [JZ23]. Distributions [BBDS17, CR17, DSS^{+}24, GPZ21]. Diversity [BCH^{+}20]. Divisible [ADH15]. Division [AD15, BK20, LV21, MBP22, Vel18]. Do [Ass20, CPS16, FG16, KR20]. Does [HM16]. Domains [BMS22, FZ15, NS15, NPJ13]. Dominance [BB17, HKV16]. Dominance-Based [BB17]. Dominant [PPS15, NPJ13]. dominant-strategy [NPJ13]. Double [AKLV17]. Dynamic [AV20, BDKS15, BV19, KT23, RSUW20]. Dynamics [AFM^{+}19, COP13, CDE^{+}14, CNP^{+}21, JLSW17, YOA^{+}13].

Each [KKL19]. Earning [BGHM19]. EC’12 [LBI15]. EC’14 [CE16]. EC’15 [FLS18]. EC’16 [CB19]. EC’17 [Ano20]. EC’19 [FM20]. EC2 [BYBYST13]. Economic [EG16]. Economics [CM13, GMMV15, GÜ23]. Edge [SSK22]. Editing [CM17]. Editorial [CM17, GÜ23, PS17]. Editors [PS17, GÜ23]. Effective [CKV20]. Effectiveness [GMS15]. Efficiency [AD15, GKK20, MBP22]. Efficient [ACV13, BCIZ21, CFGS15, CBGdK^{+}20, JT21]. Effort [KR20]. Election [ADH19]. Electric [CEK16]. Elicitation [KS19]. Eliciting [CKRS14]. Envy [AKS17, CISZ21, Elb21]. Envy-Free [AKS17]. Envy-freeness [CISZ21]. Equilibration [MBP22]. Equilibria [AEED^{+}14, BM21, BBH^{+}17, COP18, CFGS15, DS22, FS16, GKK20, GMVY18, GR16, GMC21, KYY20, KR15]. Equilibrium [BB15, GHV23, GPS13, HP19, HP21, KFH15, OPR16, RS14]. Equitable [Vel18]. Equivalence [Mau21]. ESP [JP13]. Estimates [HH13]. Evolving [AFM^{+}19]. Ex [DDW18, Mau21, RS14]. Ex-Ante [Mau21]. Ex-Post [DDW18, RS14]. Examples [BGR14]. Exchange [KMRR18, XV15]. Exclusive [KNP13]. Existence [CFGS15, GKK20]. Exploitation [GS15]. Exponential [KKM20]. Exposing [BKPS14]. Expressive [DHW15]. Extensions [PPS15]. Externalities [HIMM13]. Extraction [HH13].

Facility [AZ21, FT14]. Fair [AD15, ABM20, BK20, BCIZ21, GHLTC23, LV21]. Fairness [CKM^{+}19, CJMW20, EKPS24, MXX22, OPR16, PPS15]. Farewell [CM17]. Fast [AKLV17]. Feasible [GJLZ20, JT21]. Fee [LSZ22]. Few [AKL19]. Finding [BCIZ21, FS16]. Fine [BH22]. Fine-grained [BH22]. Finite [BBH^{+}17]. First [CKST16]. First-Price [CKST16]. Fisher [BGHM19]. Fit [SSCD14]. Fixed [GSP21, Ma22]. Fixed-Price [Ma22]. Flooded [FZ15]. Fluid [GKMP21]. Formation [BEK^{+}13]. Fractional [ABB^{+}19]. Framework [JZ23, KT23, KS19]. Free [AKS17, Elb21, KNP13]. freeness [CISZ21]. Function [MSR^{+}14]. Functions [GNR18]. Fundamental [WYZ18]. Future [CM17].

Game [ADH19, AEED^{+}14, EG16, FN13, GMC21, HKMN16, IO18, JLSW17, JZ23, JP13]. Game-Theoretic [ADH19, IO18, JZ23, GMC21, JP13]. Games [ABB^{+}19, BKPS14, BGR14, BGP15, BV19, BM21, BBH^{+}17, COP13, CFGS15, CNP^{+}21, CSV20, CG16, DS22, FS16, FTW20, FT14, GKK20, HP19, HP21, HS13, KYY20, KS23, KR15, LSCP23, MSR^{+}14, MBP22, PCFM21, PRU17, Rou15, SSK22]. Gamification [EG16]. Gap [ST20]. Generalized [FB17, GKK20, MSR^{+}14]. Geometric [FW17]. Geometry [COP13]. Given [AZ21]. Global [BKPS14]. Going [BMM18]. Good [EPR19]. Goodness [SSCD14]. Goodness-of-Fit [SSCD14]. Goods [AMS14, ADH15, BHSZ15, FSZ18, GHLTC23]. Grading [CKV20]. grained [BH22]. Graphical [COP18]. Graphs [Elb21, ST20]. Greedy [BL16]. Grid [CEK16]. Gross [GHV23]. Group [CJMW20]. Groups [AFM^{+}19].

Half [HR20]. Having [DS22]. Hedonic [ABB^{+}19]. Hereditary [LV21]. Heterogeneous [AMPS20, ADH15, DHS15, SY22]. Hierarchy [FSST19]. Homotopy [GPS13].

Housing [GHBBM22]. Howson [GPS13].

Impact [CDKS14]. Impartial [BFK17]. Imperfect [HP21, XV15]. Implementation [MSR^{+}14]. Improved [HM16]. Improving [GMS15]. Incentive [WYZ18]. Incentives [EG16, IM15, SZ20]. Incomplete [Rou15]. Inconsistent [AK21]. Increase [HM16]. Individually [BBW23]. Indivisibilities [PPS15]. Indivisible [CISZ21]. Induce [KR20]. Inefficient [LSCP23]. Inequalities [ISW23]. Inference [IO18]. Influence [ST20]. Information [BKPS14, GPTD22, KS19, Rou15]. Instance [BYBYST13]. Integrating [HLPV18]. Intensity [KLM21]. Interaction [Ass20]. Interdependent [RTC16]. Internship [BAHR18]. Introduction [Ano20, CI15, CB19, CH21a, CH21b, CE16, FN13, FLS18, FM20, HCGM22, LBI15, MT16, CM13]. Invest [KR20]. Israeli [BAHR18]. Issue [Ano20, CI15, CB19, CH21a, CH21b, CE16, FN13, FLS18, FM20, HCGM22, KKL19, LBI15]. Item [CISZ21]. Items [DHS15, Ma22].

Keyword [CBLHS15]. Knapsack [GKSA19]. Known [DSS^{+}24].

Large [AKS17, IM15, PRU17]. Leader [ADH19]. Learning [ACV13, BMM18, FP16, GMC21, JZ23, LGMC22]. Lemke [GPS13]. Leximin [KPS18]. Limit [HLPV18]. Limitations [BL16, CPV20, PPS15]. Limited [BDKS15, HN19]. Limits [BGHM19, DHS15, WYZ18]. Line [Elb21]. Linear [DGV16, GILR20, ISW23]. Liquid [GKMP21]. Liquidity [OPRS13]. Liquidity-Sensitive [OPRS13]. Lists [HPS^{+}18]. Load [VBMM22]. Local [HMV16]. Location [AZ21, FT14]. Locked [LSCP23]. Locked-in [LSCP23]. Long [MBP22]. Looking [CM17]. Lower [BH22].

Maker [OPRS13]. Makers [HLPV18]. Making [ACV13, CKRS14, MT13]. Many [BH22]. Many-to-one [BH22]. Market [ACV13, GHV23, HLPV18, LSZ22, OPRS13]. Markets [AKS17, BGHM19, CPV20, DGV16, EFRS20, HLPV18, IM15, IO18, KNP13, MXX22, Mau21, Sar13]. Marriage [CJ21]. Match [BAHR18, GMMV15]. Matching [AER^{+}24, AV20, AZ21, AKL19, ASA17, BH22, DSSX21, DSS^{+}24, FIT^{+}15, GMMV15, KYY20, MXX22, Mau21]. Matchings [CSS21, CR22]. Matroid [BCIZ21]. Matters [CNP^{+}21]. Maximin [BK20]. Maximization [BHSZ15, DDW18, DHK^{+}16, EFG^{+}14, GPTD22, MXX22, RSUW20, ST20]. Maximizers [NS15]. Maximizing [Elb21]. Maximum [ASA17, CKM^{+}19]. Measures [SSCD14]. Mechanics [GKMP21]. Mechanism [BFT16, BL16, CGMW21, DHW15, FB17, FZ15, GLT20, GK17, HH13, HMV16, KNP13, KC21, LSCP23, MS15, PT13, RTC16]. Mechanisms [ABF^{+}17, CEK16, CCK^{+}16, DSE15, FT14, GJLZ20, GHBBM22, JT21, KM24, KS19, RS14, RW18, SY22, WYZ18, WS15, NPJ13]. Mediated [Sar13]. Mediators [BFT16]. Medical [BAHR18]. Method [GPS13]. MHR [GPZ21]. Microfinance [IO18]. Minimal [FW17, GPTD22]. Minimizing [RW19]. Minimum [ASA17, FIT^{+}15]. Mining [GH24]. Mixed [ASA17]. Mixing [ABM20, EKM18]. Model [IO18]. Models [EKM18]. Modest [Ass20]. Monetary [HKV16]. Money [PT13]. Monitoring [XV15]. Monotone [JT21]. Monotonicity [RW18]. Multi [BM21, EFRS20, GMVY18]. Multi-Player [GMVY18, BM21]. Multi-Unit [EFRS20]. Multidimensional [RSUW20]. Multimarket [HV16]. Multiplayer [IO18]. Multiple [CBLHS15, KKL19, Ma22, RW19]. Multiunit [AV23]. Myopic [CDE^{+}14].

Nash [AEED^{+}14, BM21, CFGS15, CKM^{+}19, FS16, GMVY18, VBMM22]. Near [BKPS14, COP13, MS15, NW18]. Near-Optimal [MS15]. Near-Optimality [BKPS14]. Near-Optimum [NW18]. Near-Potential [COP13]. Need [Ass20]. Neighbourhood [GHBBM22]. Neighbourhood-based [GHBBM22]. Network [AEED^{+}14, BGP15, BEK^{+}13, HIMM13, RS16]. Networked [IO18]. Networks [KS23]. No [DS22]. Noisy [CPS16]. Nonequilibrium [FP16]. Nonexcludable [BHSZ15]. Nucleolus [KT23].

Observations [KNP13]. Offline

[DSSX21, MXX22]. One [SY22, BH22]. Online [ACV13, CNP^{+}21, DHK^{+}16, DSSX21, DSS^{+}24, GMMV15, KP13, MXX22, NW18, VBMM22, ZPvdS14]. Online-Matching [MXX22]. Opaque [BMM18]. Opinion [YOA^{+}13]. Opponent [GS15]. Optimal [GK16, GPZ21, GJLZ20, HIMM13, KMRR18, KM24, Ma22, MS15, NPJ13, PCFM21, RTC16, XV15]. Optimality [BKPS14]. Optimization [ACV13, DHK^{+}16, ILM^{+}21]. Optimize [CV15]. Optimized [EKPS24]. Optimum [NW18]. Oral [AKLV17]. Orders [HLPV18]. Ordinal [AZ21, CKV20, Ma22]. Our [CM17].

Packet [HPS^{+}18]. Page [DHK^{+}16]. Pairwise [CJ21]. Paradox [HV16]. parameter [NPJ13]. Pareto [KMRR18]. Part [CH21a, CH21b, HCGM22]. Participatory [GKSA19]. Partitions [GMC21]. Partners [AER^{+}24]. Payment [DFJ^{+}15]. payments [NPJ13]. Payoff [FS16]. Payoffs [EPR19]. Peer [AMPS20, CKV20, FW17]. Penalties [AK21]. Perfect [DS22, KYY20]. Performance [KR18]. Perspective [ADH19, FW17]. Pieces [AD15]. Planning [AK21]. Platforms [DSSX21, XV15]. Player [GMVY18, BM21]. Players [MS15]. Point [GSP21]. Polynomial [BV19, CG16]. Pools [GH24]. Population [LSCP23]. Position [CDE^{+}14]. Positions [AZ21, KKL19]. Positive [HIMM13]. Possibilities [CPV20]. Post [DDW18, RS14]. Posted [ABF^{+}17]. Posted-Price [ABF^{+}17]. Posting [BBDS17]. Potential [COP13]. Power [BFK17, CEK16, FT14, FZ15]. Practical [OPRS13]. Predicted [EPR19]. Prediction [AMPS20, CPV20, DSE15, FW17, HLPV18]. Predictions [CKRS14]. Preference [SSCD14]. Preferences [ABM20, BMM18, CCS17, CSS21, CR22, CJ21, Ma22]. Presence [BEK^{+}13]. Price [AER^{+}24, ABF^{+}17, AV23, BBM13, BCH^{+}20, BGR14, CDKS14, CDD^{+}14, CG16, CKST16, KS23, KM24, Ma22, MBP22, PNS16, Rou15]. Prices [BBDS17]. Pricing [AKS17, BDKS15, BYBYST13, COP18, Elb21, EFRS20, GPZ21, KKS19, RSUW20, SG20]. Principle [KC21]. Prior [KNP13]. Prior-Free [KNP13]. Priorities [BMM18, SSK22]. Priority [HPS^{+}18]. Priors [GLT20]. Privacy [CCK^{+}16, CSV20, DFI14, WYZ18]. Private [CPV20, KMRR18]. Pro [EPR19]. Pro-sociality [EPR19]. Problem [AKL19, CJ21, FB17, FP16, FSZ18]. Problems [AZ21, BCH^{+}20, BM21, BH22, DSSX21, KMM24, MS15]. Procedures [Mau21]. Profit [Elb21, HH13]. Profit-Maximizing [Elb21]. Program [DGV16]. Programming [KT23]. Projects [GNR18]. Prophet [ISW23]. Proportional [EKPS24]. Protocols [ADH19, ZPvdS14, GLR13]. Provision [CCS17]. Provision-After-Wait [CCS17]. Public [GNR18, GHBBM22, KLM21]. Pure [CFGS15, KR15]. Put [ABC^{+}15].

QPTAS [Elb21]. Quantitative [HV16].

Queries [AKL19, FS16, GMC21, LGMC22]. Query [AM14, BB15, GR16]. Question [SY22]. Quota [BCH^{+}20]. Quota-based [BCH^{+}20]. Quotas [BH22, FIT^{+}15, GHBBM22].

R [GMVY18]. R-Completeness

[GMVY18]. Random

[AMS14, FSZ18, IM15, Mau21].

Randomized [KM24]. Rank [BCIZ21]. Ranking [RGKK16]. Rating [ZPvdS14]. Ratio [KP13]. Rational [DGV16]. rationality [GLR13]. Reaching [BBW23]. Real [KMM24, KPS18]. Real-time [KMM24]. Recall [HP21]. Recaps [FP16]. Recency [FP16]. Recommendations [CKRS14]. Recommender [DFI14]. Reconstructing [BMM18]. Records [FP16]. Redesigning [BAHR18, LSZ22]. Regression [GILR20]. Regret [MBP22, RW19]. Regular [CR17, GPZ21]. Related [KMM24]. Rent [Vel18]. Repeated [DS22, PRU17]. Reserves [RW19]. Resource [PPS15]. Resources [DSSX21]. Response [CDE^{+}14, DS22, GMC21]. Restoring [KR15]. Restricted [CR22]. Results [SSCD14]. Reusable [DSSX21]. Reveal [CPS16]. Revealed [SSCD14]. Revelation [KC21]. Revenue [AKS17, BHSZ15, DDW18, EFG^{+}14, GPTD22, HM16, Ma22, RW18]. Revenue-Optimal [Ma22]. Reward [GH24, KS19]. Ride [DSSX21]. Ride-sharing [DSSX21]. Ring [CDD^{+}14]. Risk [BEK^{+}13, DSE15, PNS16]. Risk-Averse [DSE15]. Robust [GPTD22, HV16, MS15, RTC16]. Romantic [AER^{+}24]. Routing [CDD^{+}14, HPS^{+}18, MBP22, SSK22]. Rules [AM14, DFJ^{+}15, FSST19, LWC22].

Safe [GS15]. Sampling [AMS14, KP13]. Scheduling [GK17]. Schemes [EFG^{+}14, KKS19]. Scores [KR18]. Scoring [AM14, FSST19, LWC22]. Scrip [KFH15]. Search [CBLHS15, RGKK16]. Secondary [HK15]. Selection [BFK17, CJMW20, GPS13]. Selfish [CDD^{+}14, EPR19, NS15, VBMM22]. Selling [HR20, KNP13]. Sensitive [OPRS13]. Sensitivity [PNS16]. Separately [HR20]. Sequential [ABF^{+}17, AV23, BBH^{+}17, GLR13, GHBBM22, HP19, HP21, KYY20]. Service [XV15]. Services [SG20]. Set [LV21]. Shaping [CV15]. Shared [SSO23]. Shared/Buy [SSO23]. Shared/Buy-in [SSO23]. Sharing [BMS22, GKK20, GH24, RS16, DSSX21]. Shopbots [Sar13]. Shopbots-Mediated [Sar13]. Sided [CBGdK^{+}20, IM15, Mau21, DSS^{+}24]. Signaling [EFG^{+}14, FTW20, KC21, PST14]. Simple [CKV20, CGMW21, FP16, GK16, JT21, KKS19, RW18]. Simultaneous [CKST16]. Single [KKM20, WS15, NPJ13]. Single-Call [WS15]. Single-Exponential [KKM20]. single-parameter [NPJ13]. Smart [CEK16]. Social [AFM^{+}19, AZ21, PST14, SSO23, VBMM22]. sociality [EPR19]. Socially [XV15]. Socially-Optimal [XV15]. Solution [BB17, JZ23]. Solutions [GPS13]. Some [KNP13]. Sources [GILR20]. Special [Ano20, CI15, CB19, CH21a, CH21b, CE16, FLS18, FM20, HCGM22, LBI15, FN13]. Spectrum [HK15]. Spikes [EKM18]. Sponsored [CBLHS15, RGKK16]. Spot [BYBYST13]. Stability [CSS21, CG16]. Stable [BH22, BBW23, CR22, CJ21]. Stackelberg [BBH^{+}17]. Statics [HV16]. Statistical [GPTD22]. Stochastic [AKL19, EKM18, HKV16]. Strategic [BFT16, GILR20, IO18, WYZ18]. Strategically [KR20]. Strategies [DS22, SG20]. strategy [NPJ13]. Strategyproof [FIT^{+}15]. Strength [CSS21]. Strong [LGMC22]. Strongly [CR17, SY22]. Structure [CFGS15, CR22]. Structures [BBW23]. Stubborn [YOA^{+}13]. Subadditive [RW18]. Subgame [DS22, KYY20]. Subgame-Perfect [DS22]. Subjects [WYZ18]. Submodular [DHK^{+}16, FSZ18, HK15, JT21]. Subsidized [DSE15]. Substitute [GHV23]. Substitutes [LGMC22]. Supply [BDKS15]. Surrogate [LWC22]. Swap [GHBBM22]. Symmetric [BM21, GMVY18, HK15]. Systems [CEK16, DFI14, KFH15, LV21, SSO23].

Targeted [NW18]. Targeting

[HM16, HN19]. Task [DSS^{+}24]. Tasks [AV20, DSS^{+}24]. Taxes [BV19, PCFM21]. TEAC [CM17, PS17]. Team [KR18]. Technologies [LSCP23]. telling [KS19]. Terms [CM17]. Test [KR18]. Tests [SSCD14]. Theorem [GNR18, PRU17]. Theoretic [ADH19, IO18, JZ23, KS19, GMC21, JP13]. Theory [EG16, JLSW17, FN13]. Three [EPR19, SY22]. Tight [CKST16]. Tightness [BGR14]. Time [AK21, ADH15, GMS15, KKM20, KMM24]. Time-Based [GMS15]. Time-Inconsistent [AK21]. Timing [CNP^{+}21]. Topological [KS23]. Tractable [LSA24, MT13]. Trade [HLPV18]. Tradeoffs [CSS21, RGKK16]. Traffic [CV15, EKM18]. Transactions [BMM18, CM13, GÜ23]. Transfers [HKV16]. Truth [CPS16, KS19]. Truth-telling [KS19]. Truthful [ADH15, BL16, CEK16, CCK^{+}16, FB17, SY22]. Truthfulness [FZ15, HKV16]. Two [AV23, BFK17, CDD^{+}14, CBGdK^{+}20, DSS^{+}24, IM15, Mau21, SY22]. Two-Buyer [AV23]. Two-Sided [CBGdK^{+}20, IM15, Mau21, DSS^{+}24]. Types [EPR19].

Uncertainty [BBM13, PNS16]. Undirected [ST20]. Unflinchingly [EPR19]. Unified [GHLTC23]. Unit [EFRS20]. Universal [BGR14]. Unknown [BBDS17]. Unreasonable [CKM^{+}19]. Upper [BH22]. Usage [KLM21]. Users [AMPS20]. Using [EPR19]. Utilities [ASA17, GHBBM22]. Utility [BGHM19].

Valuation [DHS18]. Valuations [ABF^{+}17, BCIZ21, COP18, GJLZ20, JT21, NS15]. Value [AK21, CCK^{+}16, MSR^{+}14, WYZ18]. Values [RTC16]. VCG [DHS18, GK17, RS14]. VCG-Based [DHS18]. Verification [FZ15]. Versions [GMVY18]. Very [AKL19]. via [ACV13, FS16, GHLTC23, HH13, KT23, LGMC22, MBP22]. View [BH22]. Votes [CPS16, KKL19]. Voting [EKPS24, GKSA19, KKM20, SZ20].

Wait [CCS17]. Weak [GHV23]. Web [DHW15]. Weighted [BGR14, CFGS15, CISZ21, GKK20, KR15]. Welfare [AKS17, CKM^{+}19, DHK^{+}16, PST14, RSUW20, VBMM22]. Well [EPR19]. Whole [DHK^{+}16]. Whole-Page [DHK^{+}16]. Wild [MBP22]. WINE’13 [CI15]. WINE’18 [CH21a, CH21b]. WINE’20 [HCGM22]. without [PT13, RS16]. Workers [DSS^{+}24]. World [KPS18]. Worst [BGR14]. Worst-Case [BGR14].

References

Aziz:2019:FHG

[ABB^{+}19] Haris Aziz, Florian Brandl, Felix Brandt, Paul Harrenstein, Martin Olsen, and Dominik Peters. Fractional hedonic games. ACM Transactions on Economics and Computation, 7(2):6:1–6:??, August 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3327970.

Alon:2015:HPT

[ABC^{+}15] Noga Alon, Robert Bredereck, Jiehua Chen, Stefan Kratsch, Rolf Niedermeier, and Gerhard J. Woeginger. How to put through your agenda in collective binary decisions. ACM Transactions on Economics and Computation, 4(1): 5:1–5:??, December 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Adamczyk:2017:SPP

[ABF^{+}17] Marek Adamczyk, Allan Borodin, Diodato Ferraioli, Bart De Keijzer, and Stefano Leonardi. Sequential posted-price mechanisms with correlated valuations. ACM Transactions on Economics and Computation, 5 (4):22:1–22:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Aziz:2020:FMC

[ABM20] Haris Aziz, Anna Bogomolnaia, and Hervé Moulin. Fair mixing: The case of dichotomous preferences. ACM Transactions on Economics and Computation, 8 (4):18:1–18:27, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3417738.

Abernethy:2013:EMM

[ACV13] Jacob Abernethy, Yiling Chen, and Jennifer Wortman Vaughan. Efficient market making via convex optimization, and a connection to online learning. ACM Transactions on Economics and Computation, 1(2):12:1–12:??, May 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Aumann:2015:EFD

[AD15] Yonatan Aumann and Yair Dombb. The efficiency of fair division with connected pieces. ACM Transactions on Economics and Computation, 3 (4):23:1–23:??, July 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Aumann:2015:ATT

[ADH15] Yonatan Aumann, Yair Dombb, and Avinatan Hassidim. Auctioning time: Truthful auctions of heterogeneous divisible goods. ACM Transactions on Economics and Computation, 4(1):3:1–3:??, December 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Abraham:2019:DPL

[ADH19] Ittai Abraham, Danny Dolev, and Joseph Y. Halpern. Distributed protocols for leader election: a game-theoretic perspective. ACM Transactions on Economics and Computation, 7(1):4:1–4:??, February 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3303712.

Albers:2014:NEN

[AEED^{+}14] Susanne Albers, Stefan Eilts, Eyal Even-Dar, Yishay Mansour, and Liam Roditty. On Nash equilibria for a network creation game. ACM Transactions on Economics and Computation, 2(1):2:1–2:??, March 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Abeliuk:2024:PAA

[AER^{+}24] Andrés Abeliuk, Khaled Elbassioni, Talal Rahwan, Manuel Cebrian, and Iyad Rahwan. Price of anarchy in algorithmic matching of romantic partners. ACM Transactions on Economics and Computation, 12 (1):2:1–2:??, March 2024. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3627985.

Alon:2019:DES

[AFM^{+}19] Noga Alon, Michal Feldman, Yishay Mansour, Sigal Oren, and Moshe Tennenholtz. Dynamics of evolving social groups. ACM Transactions on Economics and Computation, 7 (3):14:1–14:??, October 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3355948.

Albers:2021:VPT

[AK21] Susanne Albers and Dennis Kraft. On the value of penalties in time-inconsistent planning. ACM Transactions on Economics and Computation, 9(3): 17:1–17:18, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3456768.

Assadi:2019:SMP

[AKL19] Sepehr Assadi, Sanjeev Khanna, and Yang Li. The stochastic matching problem with (very) few queries. ACM Transactions on Economics and Computation, 7(3):16:1–16:??, October 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3355903.

Assadi:2017:FCD

[AKLV17] Sepehr Assadi, Sanjeev Khanna, Yang Li, and Rakesh Vohra. Fast convergence in the double oral auction. ACM Transactions on Economics and Computation, 5(4):20:1–20:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Anshelevich:2017:EFP

[AKS17] Elliot Anshelevich, Koushik Kar, and Shreyas Sekar. Envyfree pricing in large markets: Approximating revenue and welfare. ACM Transactions on Economics and Computation, 5(3):16:1–16:??, August 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Azar:2014:QCS

[AM14] Pablo Daniel Azar and Silvio Micali. The query complexity of scoring rules. ACM Transactions on Economics and Computation, 2(3):10:1–10:??, July 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Agarwal:2020:PPH

[AMPS20] Arpit Agarwal, Debmalya Mandal, David C. Parkes, and Nisarg Shah. Peer prediction with heterogeneous users. ACM Transactions on Economics and Computation, 8 (1):2:1–2:34, April 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381519.

Alaei:2014:RSA

[AMS14] Saeed Alaei, Azarakhsh Malekian, and Aravind Srinivasan. On random sampling auctions for digital goods. ACM Transactions on Economics and Computation, 2(3):11:1–11:??, July 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Anonymous:2020:ISI

[Ano20] Anonymous. Introduction to the special issue on EC’17. ACM Transactions on Economics and Computation, 8 (1):1:1, April 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3383321.

Azaria:2017:DMM

[ASA17] Amos Azaria, David Sarne, and Yonatan Aumann. Distributed matching with mixed maximum–minimum utilities. ACM Transactions on Economics and Computation, 5(2): 11:1–11:??, March 2017. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Assadi:2020:CAD

[Ass20] Sepehr Assadi. Combinatorial auctions do need modest interaction. ACM Transactions on Economics and Computation, 8 (1):3:1–3:23, April 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381521.

Ahuja:2020:DMA

[AV20] Kartik Ahuja and Mihaela Van der Schaar. Dynamic matching and allocation of tasks. ACM Transactions on Economics and Computation, 7(4): 19:1–19:27, February 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369925.

Ahunbay:2023:PAT

[AV23] Mete Seref Ahunbay and Adrian Vetta. The price of anarchy of two-buyer sequential multiunit auctions. ACM Transactions on Economics and Computation, 11(1–2):2:1–2:??, June 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584864.

Anshelevich:2021:OAS

[AZ21] Elliot Anshelevich and Wennan Zhu. Ordinal approximation for social choice, matching, and facility location problems given candidate positions. ACM Transactions on Economics and Computation, 9 (2):9:1–9:24, May 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434417.

Bronfman:2018:RIM

[BAHR18] Slava Bronfman, Noga Alon, Avinatan Hassidim, and Assaf Romm. Redesigning the Israeli medical internship match. ACM Transactions on Economics and Computation, 6(3– 4):21:1–21:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3274646.

Babichenko:2015:QCC

[BB15] Yakov Babichenko and Siddharth Barman. Query complexity of correlated equilibrium. ACM Transactions on Economics and Computation, 3 (4):22:1–22:??, July 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Brandt:2017:CDB

[BB17] Felix Brandt and Markus Brill.

Computing dominance-based solution concepts. ACM Transactions on Economics and Computation, 5(2):9:1–9:??, March 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Babaioff:2017:PPU

[BBDS17] Moshe Babaioff, Liad Blumrosen, Shaddin Dughmi, and Yaron Singer. Posting prices with unknown distributions. ACM Transactions on Economics and Computation, 5(2): 13:1–13:??, March 2017. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Bosansky:2017:CSE

[BBH^{+}17] Branislav Bosanský, Simina Brânzei, Kristoffer Arnsfelt Hansen, Troels Bjerre Lund, and Peter Bro Miltersen. Computation of Stackelberg equilibria of finite sequential games. ACM Transactions on Economics and Computation, 5 (4):23:1–23:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Balcan:2013:PU

[BBM13] Maria-Florina Balcan, Avrim Blum, and Yishay Mansour. The price of uncertainty. ACM Transactions on Economics and Computation, 1(3):15:1–15:??, September 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Brandt:2023:RIS

[BBW23] Felix Brandt, Martin Bullinger, and Anaëlle Wilczynski. Reaching individually stable coalition structures. ACM Transactions on Economics and Computation, 11(1–2):4:1–4:??, June 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3588753.

Benabbou:2020:PQB

[BCH^{+}20] Nawal Benabbou, Mithun Chakraborty, Xuan-Vinh Ho, Jakub Sliwinski, and Yair Zick. The price of quota-based diversity in assignment problems. ACM Transactions on Economics and Computation, 8 (3):14:1–14:32, October 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3411513.

Benabbou:2021:FFE

[BCIZ21] Nawal Benabbou, Mithun Chakraborty, Ayumi Igarashi, and Yair Zick. Finding fair and efficient allocations for matroid rank valuations. ACM Transactions on Economics and Computation, 9(4):21:1–21:41, December 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485006.

Babaioff:2015:DPL

[BDKS15] Moshe Babaioff, Shaddin Dughmi, Robert Kleinberg, and Aleksandrs Slivkins. Dynamic pricing with limited supply. ACM Transactions on Economics and Computation, 3(1): 4:1–4:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Blume:2013:NFP

[BEK^{+}13] Lawrence Blume, David Easley, Jon Kleinberg, Robert Kleinberg, and Éva Tardos. Network formation in the presence of contagious risk. ACM Transactions on Economics and Computation, 1(2):6:1–6:??, May 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Bjelde:2017:ISP

[BFK17] Antje Bjelde, Felix Fischer, and Max Klimm. Impartial selection and the power of up to two choices. ACM Transactions on Economics and Computation, 5 (4):21:1–21:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Babaioff:2016:MDS

[BFT16] Moshe Babaioff, Moran Feldman, and Moshe Tennenholtz. Mechanism design with strategic mediators. ACM Transactions on Economics and Computation, 4(2):7:1–7:??, February 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Bei:2019:EUL

[BGHM19] Xiaohui Bei, Jugal Garg, Martin Hoefer, and Kurt Mehlhorn. Earning and utility limits in Fisher markets. ACM Transactions on Economics and Computation, 7(2):10:1–10:??, August 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3340234.

Bilo:2015:BDN

[BGP15] Davide Bilò, Luciano Gualà, and Guido Proietti. Boundeddistance network creation games. ACM Transactions on Economics and Computation, 3(3): 16:1–16:??, June 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Bhawalkar:2014:WCG

[BGR14] Kshipra Bhawalkar, Martin Gairing, and Tim Roughgarden. Weighted congestion games: The price of anarchy, universal worst-case examples, and tightness. ACM Transactions on Economics and Computation, 2(4):14:1–14:??, October 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Boehmer:2022:FGV

[BH22] Niclas Boehmer and Klaus Heeger. A fine-grained view on stable many-to-one matching problems with lower and upper quotas. ACM Transactions on Economics and Computation, 10(2):9:1–9:??, June 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546605.

Bateni:2015:RMN

[BHSZ15] Mohammadhossein Bateni, Nima Haghpanah, Balasubramanian Sivan, and Morteza Zadimoghaddam. Revenue maximization with nonexcludable goods. ACM Transactions on Economics and Computation, 3 (4):18:1–18:??, July 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Barman:2020:AAM

[BK20] Siddharth Barman and Sanath Kumar Krishnamurthy. Approximation algorithms for maximin fair division. ACM Transactions on Economics and Computation, 8(1):5:1–5:28, April 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3381525.

Balcan:2014:NOC

[BKPS14] Maria-Florina Balcan, Sara Krehbiel, Georgios Piliouras, and Jinwoo Shin. Nearoptimality in covering games by exposing global information. ACM Transactions on Economics and Computation, 2(4):13:1–13:??, October 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Borodin:2016:LGM

[BL16] Allan Borodin and Brendan Lucier. On the limitations of greedy mechanism design for truthful combinatorial auctions. ACM Transactions on Economics and Computation, 5(1):2:1–2:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Bilo:2021:CDP

[BM21] V. Bilò and M. Mavronicolas. ∃R-complete decision problems about (symmetric) Nash equilibria in (symmetric) multi-player games. ACM Transactions on Economics and Computation, 9(3):14:1–14:25, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3456758.

Blum:2018:LWG

[BMM18] Avrim Blum, Yishay Mansour, and Jamie Morgenstern. Learning what’s going on: Reconstructing preferences and priorities from opaque transactions. ACM Transactions on Economics and Computation, 6(3– 4):13:1–13:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3274642.

Birmpas:2022:CSC

[BMS22] Georgios Birmpas, Evangelos Markakis, and Guido Schäfer. Cost sharing over combinatorial domains. ACM Transactions on Economics and Computation, 10(1):4:1–4:26, March 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3505586.

Bilo:2019:DTP

[BV19] Vittorio Bilò and Cosimo Vinci.

Dynamic taxes for polynomial congestion games. ACM Transactions on Economics and Computation, 7(3):15:1– 15:??, October 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355946.

Ben-Yehuda:2013:DAE

[BYBYST13] Orna Agmon Ben-Yehuda, Muli Ben-Yehuda, Assaf Schuster, and Dan Tsafrir. Deconstructing Amazon EC2 spot instance pricing. ACM Transactions on Economics and Computation, 1 (3):16:1–16:??, September 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chen:2019:ISI

[CB19] Yiling Chen and Dirk Bergemann. Introduction to the special issue on EC’16. ACM Transactions on Economics and Computation, 7(3):11:1– 11:??, October 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3357235.

Colini-Baldeschi:2020:AET

[CBGdK^{+}20] Riccardo Colini-Baldeschi, Paul W. Goldberg, Bart de Keijzer, Stefano Leonardi, Tim Roughgarden, and Stefano Turchetta. Approximately efficient twosided combinatorial auctions. ACM Transactions on Economics and Computation, 8 (1):4:1–4:29, April 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381523.

Colini-Baldeschi:2015:MKS

[CBLHS15] Riccardo Colini-Baldeschi, Stefano Leonardi, Monika Henzinger, and Martin Starnberger. On multiple keyword sponsored search auctions with budgets. ACM Transactions on Economics and Computation, 4(1): 2:1–2:??, December 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chen:2016:TMA

[CCK^{+}16] Yiling Chen, Stephen Chong, Ian A. Kash, Tal Moran, and Salil Vadhan. Truthful mechanisms for agents that value privacy. ACM Transactions on Economics and Computation, 4 (3):13:1–13:??, June 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chan:2017:PAW

[CCS17] Hau Chan, Jing Chen, and Gowtham Srinivasan. Provisionafter-wait with common preferences. ACM Transactions on Economics and Computation, 5(2):12:1–12:??, March 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chen:2014:PAS

[CDD^{+}14] Xujin Chen, Benjamin Doerr, Carola Doerr, Xiaodong Hu, Weidong Ma, and Rob van Stee. The price of anarchy for selfish ring routing is two. ACM Transactions on Economics and Computation, 2 (2):8:1–8:??, June 2014. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Cary:2014:CPA

[CDE^{+}14] Matthew Cary, Aparna Das, Benjamin Edelman, Ioannis Giotis, Kurtis Heimerl, Anna R. Karlin, Scott Duke Kominers, Claire Mathieu, and Michael Schwarz. Convergence of position auctions under myopic best-response dynamics. ACM Transactions on Economics and Computation, 2(3):9:1– 9:??, July 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chen:2014:AIP

[CDKS14] Po-An Chen, Bart De Keijzer, David Kempe, and Guido Schäfer. Altruism and its impact on the price of anarchy. ACM Transactions on Economics and Computation, 2(4): 17:1–17:??, October 2014. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Conitzer:2016:ISI

[CE16] Vincent Conitzer and David Easley. Introduction to the special issue on EC’14. ACM Transactions on Economics and Computation, 4(4):19:1–19:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chau:2016:TMC

[CEK16] Chi-Kin Chau, Khaled Elbassioni, and Majid Khonji. Truthful mechanisms for combinatorial allocation of electric power in alternating current electric systems for smart grid. ACM Transactions on Economics and Computation, 5(1):7:1–7:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Caragiannis:2015:APN

[CFGS15] Ioannis Caragiannis, Angelo Fanelli, Nick Gravin, and Alexander Skopalik. Approximate pure Nash equilibria in weighted congestion games: Existence, efficient computation, and structure. ACM Transactions on Economics and Computation, 3(1):2:1–2:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Christodoulou:2016:PSP

[CG16] George Christodoulou and Martin Gairing. Price of stability in polynomial congestion games. ACM Transactions on Economics and Computation, 4 (2):10:1–10:??, February 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Cheng:2021:SMB

[CGMW21] Yu Cheng, Nick Gravin, Kamesh Munagala, and Kangning Wang. A simple mechanism for a budget-constrained buyer. ACM Transactions on Economics and Computation, 9(2): 10:1–10:25, May 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL

https://dl.acm.org/doi/10.
1145/3434419.

Christodoulou:2021:ISIa

[CH21a] Giorgos Christodoulou and Tobias Harks. Introduction to the special issue on WINE’18: Part 1. ACM Transactions on Economics and Computation, 9(1):1:1, March 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447510.

Christodoulou:2021:ISIb

[CH21b] Giorgos Christodoulou and Tobias Harks. Introduction to the special issue on WINE’18: Part 2. ACM Transactions on Economics and Computation, 9(2):8:1, May 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3447512.

Chen:2015:ISI

[CI15] Yiling Chen and Nicole Immorlica. Introduction to the special issue on WINE’13. ACM Transactions on Economics and Computation, 3(4):17:1–17:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chakraborty:2021:WEF

[CISZ21] Mithun Chakraborty, Ayumi Igarashi, Warut Suksompong, and Yair Zick. Weighted envyfreeness in indivisible item allocation. ACM Transactions on Economics and Computation, 9(3):18:1–18:39, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3457166.

Cseh:2021:PPS

[CJ21] Ágnes Cseh and Attila Juhos. Pairwise preferences in the stable marriage problem. ACM Transactions on Economics and Computation, 9(1):7:1– 7:28, March 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434427.

Cheng:2020:GFC

[CJMW20] Yu Cheng, Zhihao Jiang, Kamesh Munagala, and Kangning Wang. Group fairness in committee selection. ACM Transactions on Economics and Computation, 8(4):23:1–23:18, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417750.

Caragiannis:2019:UFM

[CKM^{+}19] Ioannis Caragiannis, David Kurokawa, Hervé Moulin, Ariel D. Procaccia, Nisarg Shah, and Junxing Wang. The unreasonable fairness of maximum Nash welfare. ACM Transactions on Economics and Computation, 7(3):12:1– 12:??, October 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355902.

Chen:2014:EPR

[CKRS14] Yiling Chen, Ian A. Kash, Michael Ruberry, and Victor Shnayder. Eliciting predictions and recommendations for decision making. ACM Transactions on Economics and Computation, 2(2):6:1–6:??, June 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Christodoulou:2016:TBP

[CKST16] George Christodoulou, Annamária Kovács, Alkmini Sgouritsa, and Bo Tang. Tight bounds for the price of anarchy of simultaneous first-price auctions. ACM Transactions on Economics and Computation, 4(2):9:1–9:??, February 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Caragiannis:2020:HEC

[CKV20] Ioannis Caragiannis, George A. Krimpas, and Alexandros A. Voudouris. How effective can simple ordinal peer grading be? ACM Transactions on Economics and Computation, 8 (3):16:1–16:37, October 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3412347.

Conitzer:2013:ATE

[CM13] Vincent Conitzer and R. Preston Mcafee. The ACM Transactions on Economics and Computation: an introduction. ACM Transactions on Economics and Computation, 1(1): 1:1–1:??, January 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Conitzer:2017:FEL

[CM17] Vincent Conitzer and Preston McAfee. Farewell editorial: Looking back on our terms editing ACM TEAC and into the future. ACM Transactions on Economics and Computation, 5 (2):9:1–9:??, March 2017. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Cavallo:2015:DAA

[CMV15] Ruggiero Cavallo, R. Preston Mcafee, and Sergei Vassilvitskii. Display advertising auctions with arbitrage. ACM Transactions on Economics and Computation, 3(3):15:1–15:??, June 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Chawla:2021:TMO

[CNP^{+}21] Shuchi Chawla, Joseph (Seffi)

Naor, Debmalya Panigrahi, Mohit Singh, and Seeun William Umboh. Timing matters: Online dynamics in broadcast games. ACM Transactions on Economics and Computation, 9 (2):11:1–11:22, May 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434425.

Candogan:2013:NPG

[COP13] Ozan Candogan, Asuman Ozdaglar, and Pablo A. Parrilo. Near-potential games: Geometry and dynamics. ACM Transactions on Economics and Computation, 1(2):11:1–11:??, May 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Candogan:2018:PEG

[COP18] Ozan Candogan, Asuman Ozdaglar, and Pablo Parrilo. Pricing equilibria and graphical valuations. ACM Transactions on Economics and Computation, 6(1):2:1–2:??, March 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Caragiannis:2016:WDN

[CPS16] Ioannis Caragiannis, Ariel D. Procaccia, and Nisarg Shah. When do noisy votes reveal the truth? ACM Transactions on Economics and Computation, 4 (3):15:1–15:??, June 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Cummings:2020:PLP

[CPV20] Rachel Cummings, David M. Pennock, and Jennifer Wortman Vaughan. The possibilities and limitations of private prediction markets. ACM Transactions on Economics and Computation, 8(3):15:1–15:24, October 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3412348.

Cole:2017:ASR

[CR17] Richard Cole and Shravas Rao. Applications of α-strongly regular distributions to Bayesian auctions. ACM Transactions on Economics and Computation, 5 (4):18:1–18:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Cheng:2022:SMR

[CR22] Christine T. Cheng and Will Rosenbaum. Stable matchings with restricted preferences: Structure and complexity. ACM Transactions on Economics and Computation, 10(3):13:1–13:??, September 2022. CODEN ????

ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565558.

Chen:2021:MUP

[CSS21] Jiehua Chen, Piotr Skowron, and Manuel Sorge. Matchings under preferences: Strength of stability and tradeoffs. ACM Transactions on Economics and Computation, 9(4):20:1–20:55, December 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485000.

Chen:2020:PG

[CSV20] Yiling Chen, Or Sheffet, and Salil Vadhan. Privacy games. ACM Transactions on Economics and Computation, 8 (2):9:1–9:37, May 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381533.

Chakrabarti:2015:TSO

[CV15] Deepayan Chakrabarti and Erik Vee. Traffic shaping to optimize ad delivery. ACM Transactions on Economics and Computation, 3(2):11:1–11:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Daskalakis:2018:RME

[DDW18] Constantinos Daskalakis, Nikhil R.

Devanur, and S. Matthew Weinberg. Revenue maximization and ex-post budget constraints. ACM Transactions on Economics and Computation, 6(3–4):20:1–20:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3274647.

Dandekar:2014:PAR

[DFI14] Pranav Dandekar, Nadia Fawaz, and Stratis Ioannidis. Privacy auctions for recommender systems. ACM Transactions on Economics and Computation, 2 (3):12:1–12:??, July 2014. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dutting:2015:PRT

[DFJ^{+}15] Paul Dütting, Felix Fischer, Pichayut Jirapinyo, John K. Lai, Benjamin Lubin, and David C. Parkes. Payment rules through discriminantbased classifiers. ACM Transactions on Economics and Computation, 3(1):5:1–5:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Devanur:2016:RCP

[DGV16] Nikhil R. Devanur, Jugal Garg, and László A. Végh. A rational convex program for linear Arrow–Debreu markets. ACM Transactions on Economics and Computation, 5(1): 6:1–6:??, November 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Devanur:2016:WPO

[DHK^{+}16] Nikhil R. Devanur, Zhiyi Huang, Nitish Korula, Vahab S. Mirrokni, and Qiqi Yan. Wholepage optimization and submodular welfare maximization with online bidders. ACM Transactions on Economics and Computation, 4(3):14:1–14:??, June 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dutting:2015:AHI

[DHS15] Paul Dütting, Monika Henzinger, and Martin Starnberger. Auctions for heterogeneous items and budget limits. ACM Transactions on Economics and Computation, 4(1): 4:1–4:??, December 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dutting:2018:VCV

[DHS18] Paul Dütting, Monika Henzinger, and Martin Starnberger. Valuation compressions in VCG-based combinatorial auctions. ACM Transactions on Economics and Computation, 6 (2):5:1–5:??, October 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dutting:2015:EMA

[DHW15] Paul Dütting, Monika Henzinger, and Ingmar Weber. An expressive mechanism for auctions on the Web. ACM Transactions on Economics and Computation, 4(1):1:1– 1:??, December 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dargaj:2022:DRG

[DS22] Jakub Dargaj and Jakob Grue Simonsen. Discounted repeated games having computable strategies with no computable best response under subgame-perfect equilibria. ACM Transactions on Economics and Computation, 10(1):3:1–3:39, March 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3505585.

Dimitrov:2015:SPM

[DSE15] Stanko Dimitrov, Rahul Sami, and Marina A. Epelman. Subsidized prediction mechanisms for risk-averse agents. ACM Transactions on Economics and Computation, 3(4):24:1–24:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Dickerson:2024:MTW

[DSS^{+}24] John P. Dickerson, Karthik Sankararaman, Aravind Srinivasan, Pan Xu, and Yifan Xu. Matching tasks and workers under known arrival distributions: Online task assignment with two-sided arrivals. ACM Transactions on Economics and Computation, 12(2):6:1–6:??, June 2024. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3652021.

Dickerson:2021:APR

[DSSX21] John P. Dickerson, Karthik A.

Sankararaman, Aravind Srinivasan, and Pan Xu. Allocation problems in ride-sharing platforms: Online matching with offline reusable resources. ACM Transactions on Economics and Computation, 9(3): 13:1–13:17, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3456756.

Emek:2014:SSR

[EFG^{+}14] Yuval Emek, Michal Feldman, Iftah Gamzu, Renato PaesLeme, and Moshe Tennenholtz. Signaling schemes for revenue maximization. ACM Transactions on Economics and Computation, 2(2):5:1– 5:??, June 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Ezra:2020:PMU

[EFRS20] Tomer Ezra, Michal Feldman, Tim Roughgarden, and Warut Suksompong. Pricing multi-unit markets. ACM Transactions on Economics and Computation, 7(4):20:1– 20:29, February 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3373715.

Easley:2016:IGG

[EG16] David Easley and Arpita Ghosh. Incentives, gamification, and game theory: an economic approach to badge design. ACM Transactions on Economics and Computation, 4 (3):16:1–16:??, June 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Esfandiari:2018:ATS

[EKM18] Hossein Esfandiari, Nitish Korula, and Vahab Mirrokni. Allocation with traffic spikes: Mixing adversarial and stochastic models. ACM Transactions on Economics and Computation, 6(3–4):14:1–14:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3105446.

Ebadian:2024:ODP

[EKPS24] Soroush Ebadian, Anson Kahng, Dominik Peters, and Nisarg Shah. Optimized distortion and proportional fairness in voting. ACM Transactions on Economics and Computation, 12(1):3:1–3:??, March 2024. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3640760.

Elbassioni:2021:QEF

[Elb21] Khaled Elbassioni. A QP-TAS for ϵ-envy-free profitmaximizing pricing on line graphs. ACM Transactions on Economics and Computation, 9(3):16:1–16:31, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3456762.

Epstein:2019:GBU

[EPR19] Ziv Epstein, Alexander Peysakhovich, and David Rand. The good, the bad, and the unflinchingly selfish: Pro-sociality can be well predicted using payoffs and three behavioral types. ACM Transactions on Economics and Computation, 7 (3):13:1–13:??, October 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355947.

Fadaei:2017:TMG

[FB17] Salman Fadaei and Martin Bichler. A truthful mechanism for the generalized assignment problem. ACM Transactions on Economics and Computation, 5(3):14:1–14:??, August 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Feldman:2016:DCC

[FG16] Michal Feldman and Ofir Geri.

Do capacity constraints constrain coalitions? ACM Transactions on Economics and Computation, 5(1):8:1– 8:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Fragiadakis:2015:SMM

[FIT^{+}15] Daniel Fragiadakis, Atsushi Iwasaki, Peter Troyan, Suguru Ueda, and Makoto Yokoo. Strategyproof matching with minimum quotas. ACM Transactions on Economics and Computation, 4(1):6:1–6:??, December 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Feldman:2018:ISI

[FLS18] Michal Feldman, Brendan Lucier, and Michael Schwarz. Introduction to the special issue on EC’15. ACM Transactions on Economics and Computation, 6(3–4):10:1–10:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3274531.

Fu:2020:ISI

[FM20] Hu Fu and Reshef Meir. Introduction to the special issue on EC’19. ACM Transactions on Economics and Computation, 8(4):17:1, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3424651.

Feldman:2013:ISI

[FN13] Michal Feldman and Noam Nisan. Introduction to the Special Issue on Algorithmic Game Theory. ACM Transactions on Economics and Computation, 1 (2):5:1–5:??, May 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Fudenberg:2016:RRR

[FP16] Drew Fudenberg and Alexander Peysakhovich. Recency, records, and recaps: Learning and nonequilibrium behavior in a simple decision problem. ACM Transactions on Economics and Computation, 4(4):23:1–23:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Fearnley:2016:FAN

[FS16] John Fearnley and Rahul Savani. Finding approximate Nash equilibria of bimatrix games via payoff queries. ACM Transactions on Economics and Computation, 4(4):25:1–25:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Faliszewski:2019:CSR

[FSST19] Piotr Faliszewski, Piotr Skowron, Arkadii Slinko, and Nimrod Talmon. Committee scoring rules: Axiomatic characterization and hierarchy. ACM Transactions on Economics and Computation, 7(1):3:1– 3:??, February 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3296672.

Fujishige:2018:RAP

[FSZ18] Satoru Fujishige, Yoshio Sano, and Ping Zhan. The random assignment problem with submodular constraints on goods. ACM Transactions on Economics and Computation, 6(1): 3:1–3:??, March 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Fotakis:2014:PDM

[FT14] Dimitris Fotakis and Christos Tzamos. On the power of deterministic mechanisms for facility location games. ACM Transactions on Economics and Computation, 2(4):15:1–15:??, October 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Feldman:2020:DSG

[FTW20] Moran Feldman, Moshe Tennenholtz, and Omri Weinstein. Distributed signaling games. ACM Transactions on Economics and Computation, 8 (2):7:1–7:26, May 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381529.

Frongillo:2017:GPM

[FW17] Rafael Frongillo and Jens Witkowski. A geometric perspective on minimal peer prediction. ACM Transactions on Economics and Computation, 5(3):17:1–17:??, August 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Fotakis:2015:TFD

[FZ15] Dimitris Fotakis and Emmanouil Zampetakis. Truthfulness flooded domains and the power of verification for mechanism design. ACM Transactions on Economics and Computation, 3(4):20:1–20:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Ghosh:2018:CC

[GH18] Arpita Ghosh and Patrick Hummel. Cardinal contests. ACM Transactions on Economics and Computation, 6(2): 7:1–7:??, October 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Gudmundsson:2024:BBD

[GH24] Jens Gudmundsson and Jens Leth Hougaard. Blockchain-based decentralized reward sharing: The case of mining pools. ACM Transactions on Economics and Computation, 12(1):4:1–4:??, March 2024. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3641120.

Gross-Humbert:2022:SSM

[GHBBM22] Nathanaël Gross-Humbert, Nawal Benabbou, Aurélie Beynier, and Nicolas Maudet. Sequential and swap mechanisms for public housing allocation with quotas and neighbourhood-based utilities. ACM Transactions on Economics and Computation, 10(4):15:1–15:??, December 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569704.

Gafni:2023:UFA

[GHLTC23] Yotam Gafni, Xin Huang, Ron Lavi, and Inbal Talgam-Cohen. Unified fair allocation of goods and chores via copies. ACM Transactions on Economics and Computation, 11(3–4):10:1–10:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3618116.

Garg:2023:AAM

[GHV23] Jugal Garg, Edin Husić, and László A. Végh. An auction algorithm for market equilibrium with weak gross substitute demands. ACM Transactions on Economics and Computation, 11(3–4):7:1–7:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624558.

Gast:2020:LRS

[GILR20] Nicolas Gast, Stratis Ioannidis, Patrick Loiseau, and Benjamin Roussillon. Linear regression from strategic data sources. ACM Transactions on Economics and Computation, 8 (2):10:1–10:24, May 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3391436.

Gravin:2020:OBF

[GJLZ20] Nick Gravin, Yaonan Jin, Pinyan Lu, and Chenhao Zhang. Optimal budget-feasible mechanisms for additive valuations. ACM Transactions on Economics and Computation, 8 (4):21:1–21:15, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3417746.

Ghosh:2016:OCD

[GK16] Arpita Ghosh and Robert Kleinberg. Optimal contest design for simple agents. ACM Transactions on Economics and Computation, 4(4):22:1–22:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Giannakopoulos:2017:VMB

[GK17] Yiannis Giannakopoulos and Maria Kyropoulou. The VCG mechanism for Bayesian scheduling. ACM Transactions on Economics and Computation, 5(4):19:1–19:??, December 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Gairing:2020:EEE

[GKK20] Martin Gairing, Kostas Kollias, and Grammateia Kotsialou. Existence and efficiency of equilibria for cost-sharing in generalized weighted congestion games. ACM Transactions on Economics and Computation, 8 (2):11:1–11:28, May 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3391434.

GOlz:2021:FML

[GKMP21] Paul GÖlz, Anson Kahng, Simon Mackenzie, and Ariel D. Procaccia. The fluid mechanics of liquid democracy. ACM Transactions on Economics and Computation, 9 (4):23:1–23:39, December 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485012.

Goel:2019:KVP

[GKSA19] Ashish Goel, Anilesh K. Krishnaswamy, Sukolsak Sakshuwong, and Tanja Aitamurto. Knapsack voting for participatory budgeting. ACM Transactions on Economics and Computation, 7(2):8:1–8:??, August 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3340230.

Gradwohl:2013:SRC

[GLR13] Ronen Gradwohl, Noam Livne, and Alon Rosen. Sequential rationality in cryptographic protocols. ACM Transactions on Economics and Computation, 1 (1):2:1–2:??, January 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Gergatsouli:2020:CBB

[GLT20] Evangelia Gergatsouli, Brendan Lucier, and Christos Tzamos. The complexity of black-box mechanism design with priors. ACM Transactions on Economics and Computation, 8 (4):20:1–20:19, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3417744.

Goldberg:2021:LCP

[GMC21] Paul W. Goldberg and Francisco J. Marmolejo-Cossío. Learning convex partitions and computing game-theoretic equilibria from best-response queries. ACM Transactions on Economics and Computation, 9(1):3:1–3:36, March 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434412.

Ghosh:2015:MME

[GMMV15] Arpita Ghosh, Mohammad Mahdian, R. Preston McAfee, and Sergei Vassilvitskii. To match or not to match: Economics of cookie matching in online advertising. ACM Transactions on Economics and Computation, 3(2):12:1–12:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Goldstein:2015:IET

[GMS15] Daniel G. Goldstein, R. Preston McAfee, and Siddharth Suri. Improving the effectiveness of time-based display advertising. ACM Transactions on Economics and Computation, 3(2):7:1–7:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Garg:2018:RCD

[GMVY18] Jugal Garg, Ruta Mehta, Vijay V. Vazirani, and Sadra Yazdanbod. ∃ R-completeness for decision versions of multiplayer (symmetric) Nash equilibria. ACM Transactions on Economics and Computation, 6 (1):1:1–1:??, March 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Gopalan:2018:PPB

[GNR18] Parikshit Gopalan, Noam Nisan, and Tim Roughgarden. Public projects, Boolean functions, and the borders of Border’s theorem. ACM Transactions on Economics and Computation, 6(3–4):18:1– 18:??, November 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3274645.

Goldberg:2013:CHM

[GPS13] Paul W. Goldberg, Christos H. Papadimitriou, and Rahul Savani. The complexity of the homotopy method, equilibrium selection, and Lemke–Howson solutions. ACM Transactions on Economics and Computation, 1 (2):9:1–9:??, May 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Giannakopoulos:2022:RRM

[GPTD22] Yiannis Giannakopoulos, Diogo Poças, and Alexandros Tsigonias-Dimitriadis. Robust revenue maximization under minimal statistical information. ACM Transactions on Economics and Computation, 10(3):11:1–11:??, September 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3546606.

Giannakopoulos:2021:OPM

[GPZ21] Yiannis Giannakopoulos, Diogo Poças, and Keyu Zhu. Optimal pricing for MHR and λ-regular distributions. ACM Transactions on Economics and Computation, 9(1):2:1–2:28, March 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434423.

Goldberg:2016:BQC

[GR16] Paul W. Goldberg and Aaron Roth. Bounds for the query complexity of approximate equilibria. ACM Transactions on Economics and Computation, 4(4):24:1–24:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Ganzfried:2015:SOE

[GS15] Sam Ganzfried and Tuomas Sandholm. Safe opponent exploitation. ACM Transactions on Economics and Computation, 3(2):8:1–8:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Ganor:2021:CCF

[GSP21] Anat Ganor, Karthik C. S., and Dömötör Pálvölgyi. On communication complexity of fixed point computation. ACM Transactions on Economics and Computation, 9(4):25:1–25:27, December 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3485004.

Goldberg:2023:ENC

[GÜ23] Paul Goldberg and Utku Ünver. Editorial from the new co-Editors-in-Chief of ACM Transactions on Economics and Computation. ACM Transactions on Economics and Computation, 11(3–4):6:1–6:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631669.

Hoefer:2022:ISI

[HCGM22] Martin Hoefer, Xujin Chen, Nikolai Gravin, and Ruta Mehta. Introduction to the special issue on WINE’20: Part 1. ACM Transactions on Economics and Computation, 10 (2):6:1–6:??, June 2022. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3555339.

Ha:2013:MDC

[HH13] Bach Q. Ha and Jason D.

Hartline. Mechanism design via consensus estimates, cross checking, and profit extraction. ACM Transactions on Economics and Computation, 1 (2):8:1–8:??, May 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Haghpanah:2013:OAP

[HIMM13] Nima Haghpanah, Nicole Immorlica, Vahab Mirrokni, and Kamesh Munagala. Optimal auctions with positive network externalities. ACM Transactions on Economics and Computation, 1(2):13:1–13:??, May 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hoefer:2015:SSA

[HK15] Martin Hoefer and Thomas Kesselheim. Secondary spectrum auctions for symmetric and submodular bidders. ACM Transactions on Economics and Computation, 3(2): 9:1–9:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hassidim:2016:G

[HKMN16] Avinatan Hassidim, Haim Kaplan, Yishay Mansour, and Noam Nisan. The AND–OR game. ACM Transactions on Economics and Computation, 5(1):1:1–1:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hoefer:2016:TSD

[HKV16] Martin Hoefer, Thomas Kesselheim, and Berthold Vöcking. Truthfulness and stochastic dominance with monetary transfers. ACM Transactions on Economics and Computation, 4(2):11:1–11:??, February 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Heidari:2018:IMM

[HLPV18] Hoda Heidari, Sébastien Lahaie, David M. Pennock, and Jennifer Wortman Vaughan. Integrating market makers, limit orders, and continuous trade in prediction markets. ACM Transactions on Economics and Computation, 6(3–4):15:1– 15:??, November 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3274643.

Hummel:2016:WDI

[HM16] Patrick Hummel and R. Preston Mcafee. When does improved targeting increase revenue? ACM Transactions on Economics and Computation, 5(1):4:1–4:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hassidim:2016:LCM

[HMV16] Avinatan Hassidim, Yishay Mansour, and Shai Vardi. Local computation mechanism design. ACM Transactions on Economics and Computation, 4(4):21:1–21:??, August 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hummel:2019:BLT

[HN19] Patrick Hummel and Uri Nadav. Bid-limited targeting. ACM Transactions on Economics and Computation, 7(2): 5:1–5:??, August 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3327968.

Halpern:2019:SEC

[HP19] Joseph Y. Halpern and Rafael Pass. Sequential equilibrium in computational games. ACM Transactions on Economics and Computation, 7(2): 9:1–9:??, August 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340232.

Halpern:2021:SEG

[HP21] Joseph Y. Halpern and Rafael Pass. Sequential equilibrium in games of imperfect recall. ACM Transactions on Economics and Computation, 9 (4):22:1–22:26, December 2021.

CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3485002.

Harks:2018:CPR

[HPS^{+}18] Tobias Harks, Britta Peis, Daniel Schmand, Bjoern Tauer, and Laura Vargas Koch. Competitive packet routing with priority lists. ACM Transactions on Economics and Computation, 6(1):4:1–4:??, March 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Hart:2020:BHS

[HR20] Sergiu Hart and Philip J. Reny. The better half of selling separately. ACM Transactions on Economics and Computation, 7 (4):18:1–18:18, February 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3369927.

Hoefer:2013:AAC

[HS13] Martin Hoefer and Alexander Skopalik. Altruism in atomic congestion games. ACM Transactions on Economics and Computation, 1(4):21:1–21:??, December 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Harks:2016:RQC

[HV16] Tobias Harks and Philipp Von Falkenhausen. Robust quantitative comparative statics for a multimarket paradox. ACM Transactions on Economics and Computation, 5(1):3:1– 3:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Immorlica:2021:CAO

[ILM^{+}21] Nicole Immorlica, Brendan Lucier, Jieming Mao, Vasilis Syrgkanis, and Christos Tzamos. Combinatorial assortment optimization. ACM Transactions on Economics and Computation, 9(1):5:1–5:34, March 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434415.

Immorlica:2015:ILR

[IM15] Nicole Immorlica and Mohammad Mahdian. Incentives in large random two-sided markets. ACM Transactions on Economics and Computation, 3 (3):14:1–14:??, June 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Irfan:2018:CSI

[IO18] Mohammad T. Irfan and Luis E. Ortiz. Causal strategic inference in a game-theoretic model of multiplayer networked microfinance markets. ACM Transactions on Economics and Computation, 6(2):6:1– 6:??, October 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Immorlica:2023:PIL

[ISW23] Nicole Immorlica, Sahil Singla, and Bo Waggoner. Prophet inequalities with linear correlations and augmentations. ACM Transactions on Economics and Computation, 11 (3–4):8:1–8:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3623273.

Jaggard:2017:DBG

[JLSW17] Aaron D. Jaggard, Neil Lutz, Michael Schapira, and Rebecca N. Wright. Dynamics at the boundary of game theory and distributed computing. ACM Transactions on Economics and Computation, 5(3): 15:1–15:??, August 2017. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Jain:2013:GTA

[JP13] Shaili Jain and David C. Parkes. A game-theoretic analysis of the ESP game. ACM Transactions on Economics and Computation, 1(1): 3:1–3:??, January 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Jalaly:2021:SEB

[JT21] Pooya Jalaly and Éva Tardos. Simple and efficient budget feasible mechanisms for monotone submodular valuations. ACM Transactions on Economics and Computation, 9 (1):4:1–4:20, March 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434421.

Jha:2023:LFD

[JZ23] Tushant Jha and Yair Zick.

A learning framework for distribution-based game-theoretic solution concepts. ACM Transactions on Economics and Computation, 11(1–2):5:1–5:??, June 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580374.

Kephart:2021:RPM

[KC21] Andrew Kephart and Vincent Conitzer. The revelation principle for mechanism design with signaling costs. ACM Transactions on Economics and Computation, 9(1):6:1–6:35, March 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434408.

Kash:2015:EAS

[KFH15] Ian A. Kash, Eric J. Friedman, and Joseph Y. Halpern. An equilibrium analysis of scrip systems. ACM Transactions on Economics and Computation, 3 (3):13:1–13:??, June 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Kirousis:2019:AVM

[KKL19] Lefteris Kirousis, Phokion G.

Kolaitis, and John Livieratos. Aggregation of votes with multiple positions on each issue.

ACM Transactions on Economics and Computation, 7 (1):1:1–1:??, February 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3296675.

Knop:2020:VBS

[KKM20] Dusan Knop, Martin Koutecký, and Matthias Mnich. Voting and bribing in singleexponential time. ACM Transactions on Economics and Computation, 8(3):12:1–12:28, October 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3396855.

Kash:2019:SPS

[KKS19] Ian A. Kash, Peter Key, and Warut Suksompong. Simple pricing schemes for the cloud. ACM Transactions on Economics and Computation, 7(2): 7:1–7:??, August 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3327973.

Kannan:2021:CUI

[KLM21] Aadharsh Kannan, Jacob LaRiviere, and R. Preston McAfee. Characterizing the usage intensity of public cloud. ACM Transactions on Economics and Computation, 9(3):15:1–15:18, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3456760.

Ko:2024:OPD

[KM24] Shao-Heng Ko and Kamesh Munagala. Optimal price discrimination for randomized mechanisms. ACM Transactions on Economics and Computation, 12(2):5:1–5:??, June 2024. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3650107.

Kinnear:2024:CRT

[KMM24] Ryan Kinnear, Ravi R. Mazumdar, and Peter Marbach. Convexity in real-time bidding and related problems. ACM Transactions on Economics and Computation, 12(2):7:1–7:??, June 2024. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656552.

Kannan:2018:PPO

[KMRR18] Sampath Kannan, Jamie Morgenstern, Ryan Rogers, and Aaron Roth. Private Pareto optimal exchange. ACM Transactions on Economics and Computation, 6(3–4):12:1– 12:??, November 2018. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3105445.

Karlin:2013:SEM

[KNP13] Anna R. Karlin, C. Thach Nguyen, and Yuval Peres. Selling in exclusive markets: Some observations on prior-free mechanism design. ACM Transactions on Economics and Computation, 1(2):7:1–7:??, May 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Koutsoupias:2013:CRO

[KP13] Elias Koutsoupias and George Pierrakos. On the competitive ratio of online sampling auctions. ACM Transactions on Economics and Computation, 1 (2):10:1–10:??, May 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Kurokawa:2018:LAR

[KPS18] David Kurokawa, Ariel D. Procaccia, and Nisarg Shah. Leximin allocations in the real world. ACM Transactions on Economics and Computation, 6(3–4):11:1–11:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?id=
                                        3274641.

Kollias:2015:RPE

[KR15] Konstantinos Kollias and Tim Roughgarden. Restoring pure equilibria to weighted congestion games. ACM Transactions on Economics and Computation, 3(4):21:1–21:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Kleinberg:2018:TPT

[KR18] Jon Kleinberg and Maithra Raghu. Team performance with test scores. ACM Transactions on Economics and Computation, 6(3–4):17:1–17:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3274644.

Kleinberg:2020:HDC

[KR20] Jon Kleinberg and Manish Raghavan. How do classifiers induce agents to invest effort strategically? ACM Transactions on Economics and Computation, 8(4):19:1–19:23, November 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417742.

Kong:2019:ITF

[KS19] Yuqing Kong and Grant Schoenebeck. An information theoretic framework for designing information elicitation mechanisms that reward truthtelling. ACM Transactions on Economics and Computation, 7(1):2:1–2:??, February 2019. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3296670.

Kleer:2023:TBP

[KS23] Pieter Kleer and Guido Schäfer. Topological bounds on the price of anarchy of clustering games on networks. ACM Transactions on Economics and Computation, 11(3–4):11:1–11:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3625689.

Koenemann:2023:FCN

[KT23] Jochen Koenemann and Justin Toth. A framework for computing the nucleolus via dynamic programming. ACM Transactions on Economics and Computation, 11(1–2):3:1–3:??, June 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580375.

Kawase:2020:SPE

[KYY20] Yasushi Kawase, Yutaro Yamaguchi, and Yu Yokoi. Subgame perfect equilibria of sequential matching games. ACM Transactions on Economics and Computation, 7(4):21:1– 21:30, February 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3373717.

Leyton-Brown:2015:ISI

[LBI15] Kevin Leyton-Brown and Panos Ipeirotis. Introduction to the special issue on EC’12.

ACM Transactions on Economics and Computation, 3(1): 1:1–1:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Lock:2022:LSS

[LGMC22] Edwin Lock, Paul W. Goldberg, and Francisco Marmolejo-Cossío. Learning strong substitutes demand via queries. ACM Transactions on Economics and Computation, 10 (2):10:1–10:??, June 2022. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546604.

Levy:2024:TBC

[LSA24] Priel Levy, David Sarne, and Yonatan Aumann. Tractable binary contests. ACM Transactions on Economics and Computation, 12(1):1:1–1:??, March 2024. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3630109.

Leonardos:2023:CDP

[LSCP23] Stefanos Leonardos, Joseph Sakos, Costas Courcoubetis, and Georgios Piliouras. Catastrophe by design in population games: a mechanism to destabilize inefficient locked-in technologies. ACM Transactions on Economics and Computation, 11(1–2):1:1–1:??, June 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583782.

Lavi:2022:RBF

[LSZ22] Ron Lavi, Or Sattath, and Aviv Zohar. Redesigning Bitcoin’s fee market. ACM Transactions on Economics and Computation, 10(1):5:1–5:31, March 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3530799.

Li:2021:FDH

[LV21] Z. Li and A. Vetta. The fair division of hereditary set systems. ACM Transactions on Economics and Computation, 9 (2):12:1–12:19, May 2021. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3434410.

Liu:2022:SSR

[LWC22] Yang Liu, Juntao Wang, and Yiling Chen. Surrogate scoring rules. ACM Transactions on Economics and Computation, 10(3):12:1–12:??, September 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3565559.

Ma:2022:ROD

[Ma22] Will Ma. Revenue-optimal deterministic auctions for multiple buyers with ordinal preferences over fixed-price items. ACM Transactions on Economics and Computation, 10 (2):7:1–7:??, June 2022. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3555045.

Mauras:2021:TSR

[Mau21] Simon Mauras. Two-sided random matching markets: Exante equivalence of the deferred acceptance procedures. ACM Transactions on Economics and Computation, 9 (4):24:1–24:14, December 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485010.

Monnot:2022:RGW

[MBP22] Barnabé Monnot, Francisco Benita, and Georgios Piliouras. Routing games in the wild: Efficiency, equilibration, regret, and a price of anarchy bound via long division. ACM Transactions on Economics and Computation, 10(1):2:1–2:26, March 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3512747.

Minooei:2015:NOR

[MS15] Hadi Minooei and Chaitanya Swamy. Near-optimal and robust mechanism design for covering problems with correlated players. ACM Transactions on Economics and Computation, 3 (4):19:1–19:??, July 2015. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Michalak:2014:ICV

[MSR^{+}14] Tomasz P. Michalak, Piotr L. Szczepański, Talal Rahwan, Agata Chrobak, Simina Brânzei, Michael Wooldridge, and Nicholas R. Jennings. Implementation and computation of a value for generalized characteristic function games. ACM Transactions on Economics and Computation, 2(4):16:1–16:??, October 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Mossel:2013:MCT

[MT13] Elchanan Mossel and Omer Tamuz. Making consensus tractable. ACM Transactions on Economics and Computation, 1(4):20:1–20:??, December 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Mcafee:2016:I

[MT16] Preston Mcafee and Éva Tardos. Introduction. ACM Transactions on Economics and Computation, 4(3):12:1–12:??, June 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Ma:2022:FMA

[MXX22] Will Ma, Pan Xu, and Yifan Xu. Fairness maximization among offline agents in online-matching markets. ACM Transactions on Economics and Computation, 10(4):16:1–16:??, December 2022. CODEN ????

ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569705.

Naroditskiy:2013:OPD

[NPJ13] Victor Naroditskiy, Maria Polukarov, and Nicholas R. Jennings. Optimal payments in dominant-strategy mechanisms for single-parameter domains. ACM Transactions on Economics and Computation, 1 (1):4:1–4:??, January 2013. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Nath:2015:AMD

[NS15] Swaprava Nath and Arunava Sen. Affine maximizers in domains with selfish valuations. ACM Transactions on Economics and Computation, 3(4): 26:1–26:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Naor:2018:NOO

[NW18] Joseph (Seffi) Naor and David Wajc. Near-optimum online ad allocation for targeted advertising. ACM Transactions on Economics and Computation, 6(3–4):16:1–16:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3105447.

Othman:2016:CFT

[OPR16] Abraham Othman, Christos Papadimitriou, and Aviad Rubinstein. The complexity of fairness through equilibrium. ACM Transactions on Economics and Computation, 4(4): 20:1–20:??, August 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Othman:2013:PLS

[OPRS13] Abraham Othman, David M. Pennock, Daniel M. Reeves, and Tuomas Sandholm. A practical liquidity-sensitive automated market maker. ACM Transactions on Economics and Computation, 1(3):14:1–14:??, September 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Paccagnan:2021:OTA

[PCFM21] Dario Paccagnan, Rahul Chandan, Bryce L. Ferguson, and Jason R. Marden. Optimal taxes in atomic congestion games. ACM Transactions on Economics and Computation, 9(3): 19:1–19:33, September 2021. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457168.

Piliouras:2016:RSP

[PNS16] Georgios Piliouras, Evdokia Nikolova, and Jeff S. Shamma. Risk sensitivity of price of anarchy under uncertainty. ACM Transactions on Economics and Computation, 5(1):5:1– 5:??, November 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Parkes:2015:BDR

[PPS15] David C. Parkes, Ariel D. Procaccia, and Nisarg Shah. Beyond dominant resource fairness: Extensions, limitations, and indivisibilities. ACM Transactions on Economics and Computation, 3(1):3:1– 3:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Pai:2017:ATL

[PRU17] Mallesh M. Pai, Aaron Roth, and Jonathan Ullman. An antifolk theorem for large repeated games. ACM Transactions on Economics and Computation, 5(2):10:1–10:??, March 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Pennock:2017:ENT

[PS17] David Pennock and Ilya Segal.

Editorial from the new TEAC Co-Editors-in-Chief. ACM Transactions on Economics and Computation, 5(2):9:1– 9:??, March 2017. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Polevoy:2014:SCS

[PST14] Gleb Polevoy, Rann Smorodinsky, and Moshe Tennenholtz. Signaling competition and social welfare. ACM Transactions on Economics and Computation, 2(1):1:1–1:??, March 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Procaccia:2013:AMD

[PT13] Ariel D. Procaccia and Moshe Tennenholtz. Approximate mechanism design without money. ACM Transactions on Economics and Computation, 1 (4):18:1–18:??, December 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Roberts:2016:RTS

[RGKK16] Ben Roberts, Dinan Gunawardena, Ian A. Kash, and Peter Key. Ranking and tradeoffs in sponsored search auctions. ACM Transactions on Economics and Computation, 4 (3):17:1–17:??, June 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Roughgarden:2015:PAG

[Rou15] Tim Roughgarden. The price of anarchy in games of incomplete information. ACM Transactions on Economics and Computation, 3(1):6:1–6:??, March 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Rozen:2014:EPE

[RS14] Rakefet Rozen and Rann Smorodinsky. Ex-post equilibrium and VCG mechanisms. ACM Transactions on Economics and Computation, 2(2): 7:1–7:??, June 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Roughgarden:2016:NCS

[RS16] Tim Roughgarden and Okke Schrijvers. Network costsharing without anonymity. ACM Transactions on Economics and Computation, 4(2): 8:1–8:??, February 2016. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Roth:2020:MDP

[RSUW20] Aaron Roth, Aleksandrs Slivkins, Jonathan Ullman, and Zhiwei Steven Wu. Multidimensional dynamic pricing for welfare maximization. ACM Transactions on Economics and Computation, 8(1):6:1– 6:35, April 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381527.

Roughgarden:2016:ORM

[RTC16] Tim Roughgarden and Inbal Talgam-Cohen. Optimal and robust mechanism design with interdependent values. ACM Transactions on Economics and Computation, 4(3):18:1–18:??, June 2016. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Rubinstein:2018:SMS

[RW18] Aviad Rubinstein and S. Matthew Weinberg. Simple mechanisms for a subadditive buyer and applications to revenue monotonicity. ACM Transactions on Economics and Computation, 6(3–4):19:1–19:??, November 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3105448.

Roughgarden:2019:MRM

[RW19] Tim Roughgarden and Joshua R. Wang. Minimizing regret with multiple reserves. ACM Transactions on Economics and Computation, 7(3):17:1– 17:??, October 2019. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355900.

Sarne:2013:CSM

[Sar13]  David Sarne.  Competitive

shopbots-mediated markets. ACM Transactions on Economics and Computation, 1 (3):17:1–17:??, September 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Song:2020:PBS

[SG20] Jiayi Song and Roch Guérin. Pricing (and bidding) strategies for delay differentiated cloud services. ACM Transactions on Economics and Computation, 8 (2):8:1–8:58, May 2020. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381531.

Smeulders:2014:GFM

[SSCD14] Bart Smeulders, Frits C. R. Spieksma, Laurens Cherchye, and Bram De Rock. Goodnessof-fit measures for revealed preference tests: Complexity results and algorithms. ACM Transactions on Economics and Computation, 2(1):3:1–3:??, March 2014. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Scheffler:2022:RGE

[SSK22] Robert Scheffler, Martin Strehler, and Laura Vargas Koch. Routing games with edge priorities. ACM Transactions on Economics and Computation, 10(1):1:1–1:27, March 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3488268.

Shi:2023:SCA

[SSO23] Zhenpeng Shi, David Starobinski, and Ariel Orda. Social cost analysis of shared/buy-in computing systems. ACM Transactions on Economics and Computation, 11(3–4):9:1–9:??, December 2023. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624355.

Schoenebeck:2020:IMU

[ST20] Grant Schoenebeck and Biaoshuai Tao. Influence maximization on undirected graphs: Toward closing the (1–1/e) gap. ACM Transactions on Economics and Computation, 8(4):22:1–22:36, November 2020. CODEN ????

ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3417748.

Schoenebeck:2022:TST

[SY22] Grant Schoenebeck and Fang-Yi Yu. Two strongly truthful mechanisms for three heterogeneous agents answering one question. ACM Transactions on Economics and Computation, 10(4):14:1–14:??, December 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3565560.

Shah:2020:AVI

[SZ20] Nihar B. Shah and Dengyong Zhou. Approval voting and incentives in crowdsourcing. ACM Transactions on Economics and Computation, 8 (3):13:1–13:40, October 2020. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3396863.

Vinci:2022:NSW

[VBMM22] Cosimo Vinci, Vittorio Bilò, Gianpiero Monaco, and Luca Moscardelli. Nash social welfare in selfish and online load balancing. ACM Transactions on Economics and Computation, 10(2):8:1–8:??, June 2022. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3544978.

Velez:2018:ERD

[Vel18] Rodrigo A. Velez. Equitable rent division. ACM Transactions on Economics and Computation, 6(2):9:1–9:??, October 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Wilkens:2015:SCM

[WS15] Christopher A. Wilkens and Balasubramanian Sivan. Singlecall mechanisms. ACM Transactions on Economics and Computation, 3(2):10:1–10:??, April 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Wang:2018:VPS

[WYZ18] Weina Wang, Lei Ying, and Junshan Zhang. The value of privacy: Strategic data subjects, incentive mechanisms, and fundamental limits. ACM Transactions on Economics and Computation, 6(2):8:1– 8:??, October 2018. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Xiao:2015:SOD

[XV15] Yuanzhang Xiao and Mihaela Van Der Schaar. Sociallyoptimal design of service exchange platforms with imperfect monitoring. ACM Transactions on Economics and Computation, 3(4):25:1–25:??, July 2015. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Yildiz:2013:BOD

[YOA^{+}13] Ercan Yildiz, Asuman Ozdaglar, Daron Acemoglu, Amin Saberi, and Anna Scaglione. Binary opinion dynamics with stubborn agents. ACM Transactions on Economics and Computation, 1(4):19:1–19:??, December 2013. CODEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).

Zhang:2014:RPO

[ZPvdS14] Yu Zhang, Jaeok Park, and Mihaela van der Schaar. Rating protocols in online communities. ACM Transactions on Economics and Computation, 2 (1):4:1–4:??, March 2014. CO-DEN ???? ISSN 2167-8375 (print), 2167-8383 (electronic).